how do i solve by substitution -x+2y=17
Answer by ilana(307)

You can't. Since you have 2 variables (x and y), you need 2 equations in order to solve by substitution (or to solve the system in general). If you had another equation, such as x+y=1, you could solve by substitution. So I will solve the system I just made up of -x+2y=17 and x+y=1. Substitution means solving one equation for x or y, and then plugging that new "x" or "y" into the other equation. So for my system, I could take the second equation, x+y=1, and solve for x: x=1-y. Now anywhere I have an x, I can put 1-y. So the first equation, -x+2y=17, becomes -(1-y)+2y=17. Distribute the -1 to get -1+y+2y=17. Simplify to get -1+3y=17. Add 1 to both sides: 3y=18. Divide both sides by 3: y=6. Now plug that value of y back into either of the equations to solve for x. x+y=1, so x+(6)=1, so x=-5. So the solution to this system is x=-5, y=6. Those values work in both equations in the system.